KittyGirl:

Simplify negative 2 and 1 over 6 – negative 7 and 1 over 3. negative 5 and 1 over 6 negative 9 and 3 over 6 9 and 3 over 6 5 and 1 over 6

kittybasil:

What I got from this:

Quote:
Simplify \(-2\) and \(\frac{1}{6}-(-7)\) and \(\frac{1}{3}\)
Is this what you meant?
Your given function would translate to:\[-2+\frac{1}{6}-(-7)+\frac{1}{3}\]

kittybasil:

Step one: make everything denominator 6. The red parts are what help you do this.\[-2(\color{red}{\frac{6}{6}})+\frac{1}{6}-(-7)(\color{red}{\frac{6}{6}})+\frac{1}{3}(\color{red}{\frac{2}{2}})\]

kittybasil:

Step two: simplify by multiplication (anything without a denominator automatically has 1 at the bottom) and you get:\[\frac{-2}{6}+\frac{1}{6}-(\frac{-42}{6})+\frac{2}{6}\]

kittybasil:

Step three: add the numerators, and keep the denominator.\[\frac{-2+1-(-42)+2}{6}\]\[\frac{1-(-42)}{6}=\frac{43}{6}\]

Shadow:

I think they mean: \[-2 \frac{ 1 }{ 6 } - - 7 \frac{ 1 }{ 3 }\] \[-2 \frac{ 1 }{ 6 } + 7 \frac{ 1 }{ 3 }\]

kittybasil:

Aw

Shadow:

\[-2 = - \frac{ 12 }{ 6 }\] So, \[- \frac{ 12 }{ 6 } + - \frac{ 1 }{ 6} = - \frac{ 13 }{ 6 }\] \[7 = \frac{ 21 }{ 3 }\] So, \[\frac{ 21 }{ 3 } + \frac{ 1 }{ 3} = \frac{ 22 }{ 3 }\] You can only add fractions that have the same denominator so, \[\frac{ 2 }{ 2 } \times \frac{ 22 }{ 3 } = \frac{ 44 }{ 6}\] Thus our problem is, \[-\frac{ 13 }{ 6 } + \frac{ 44 }{ 6 } = \frac{ 31 }{ 6 }\]